The concentration of N2O5 is initially 0.500 mol/L. After 100 s it has dropped to 0.250 mol/L and after 200 s it has dropped to 0.125 mol/L. What is the concentration after 150 s?

Use the first order concentration-time equation to determine [N2O5] remaining after 150 s.
Answer #1

The decomposition of N2O5 is a first order reaction.

Given that Initial concentration of N2O5, [N2O5]0 = 0.500 mol/L

Concentration of N2O5 after 100s become [N2O5]100 = 0.250 mol/L

i.e. after 100s initial concentration become half, so the half life period is 100s

= 0.693/K

K = 0.693 /t12 = = = 0.00693 S-1

We have the first order equation as

K = 2.303150 log 0.500[A]

0.00693 = 2.303150 log 0.500[A]

0.4513 = log (0.500) – log [A]

Log [A] = -0.7523

[A]150 = 0.176 mol/L